OpenStudy (anonymous):

x^2/(x^2-3) - (x+1)/(x+3) = 0, solve for x.....any help?

OpenStudy (dumbcow):

\[\frac{x^2}{x^{2}-3} - \frac{x+1}{x+3} = 0\] \[\frac{x^2}{x^{2}-3} = \frac{x+1}{x+3}\] cross-multiply \[x^{2}(x+3) = (x+1)(x^{2}-3)\] \[x^{3} +3x^{2} = x^{3}+x^{2}-3x-3\] Combine like terms \[2x^{2}+3x+3 = 0\] Use quadratic formula \[x = \frac{-3 \pm \sqrt{3^{2}-4(2)(3)}}{2(2)}\] \[x = \frac{-3 \pm \sqrt{15}i}{4}\] No real solution